1. Determine whether the set of all vectors of the form (a, b ,c) such that b=a+c+1 is a subspace of R3 .

2. Determine whether the vector p is in the span {S}. Given S = { p_1,p_2,p_3} where
p_1=2+x+〖4x〗^2, p_2=1-x+〖3x〗^2, p_3=3+2x+〖5x〗^2; p=7+8x+9x^2.

3. Using two methods, verify that the set S is linearly independent.
v1 = (2, -2, 0), v2=(6,1,4), v3=(2,0,-4)

4. Express v as a linear combination of the vectors v_(1,), v_2,v_3.

v=(2,-1,3), v1=(1,0,0), v2=(2,2,0), v3=(3,3,3)
1
Expert's answer
2019-12-16T11:17:37-0500
  1. For "V_1 = (a, a+c+1, c), a \in \R, c \in \R" to be a subspace of "\R^3" , it must be a vector space itself. Let us check the axioms: "\bold v_1 + \bold v_2 = (a_1, a_1 + c_1 + 1, c_1) + (a_2, a_2 + c_2 + 1, c_2)" "\bold v_1 + \bold v_2 = (a_1 + a_2, (a_1+a_2) + (c_1+c_2) + 2, c_1 + c_2) \notin V_1", hence "V_1" is not a subspace.
  2. For "\bold p \in span(\bold p_1, \bold p_2, \bold p_3)", there must exist unique "\lambda_1, \lambda_2, \lambda_3", not all equal to zero, so that "\bold p = \lambda_1 \bold p_1 + \lambda_2 \bold p_2 + \lambda_3 \bold p_3". Writing the last equation in vector form, obtain: "\left(\begin{array}{ccc|c} 
 2 & 1 & 3 & 7\\ 
 1 & -1 & 2 & 8\\
16 & 9 & 25 & 9 
\end{array}\right)". Performing Gauss-Jordan elimination on last matrix, obtain "\left(\begin{array}{ccc|c} 
 1 & 0 & 0 & 60\\ 
 0 & 1 & 0 & -14\\
 0 & 0 & 1 & -33 
\end{array}\right)", hence "\lambda_1 = 60, \lambda_2 = -14, \lambda_3 = -33" and vector "\bold p" is in span of "\bold p_1, \bold p_2, \bold p_3".
  3. First way: "\bold v_1, \bold v_2, \bold v_3" are linearly independent, if "\lambda_1 \bold v_1 + \lambda_2 \bold v_2 + \lambda_3 \bold v_3 = 0" only when all "\lambda_i = 0". Rewriting the last equation in vector form, obtain "\left(\begin{array}{ccc}
 2 & 6 & 2\\ 
 -2 & 1 & 0\\
0 & 4 & -4 
\end{array}\right)\left(\begin{array}{c}
 \lambda_1\\
\lambda_2\\
\lambda_3 
\end{array}\right) = \bold 0". Reducing the matrix, obtain "\left(\begin{array}{ccc}
 1 & 0 & 0\\ 
 0 & 1 & 0\\
0 & 0 & 1 
\end{array}\right)\left(\begin{array}{c}
 \lambda_1\\
\lambda_2\\
\lambda_3 
\end{array}\right) = \bold 0", from where all "\lambda_i = 0", hence "\bold v_1, \bold v_2, \bold v_3" are linearly independent. Second way: Vectors are linearly independent, if the determinant of matrix, spanned by vectors is zero. "det \left(\begin{array}{ccc}
 2 & 6 & 2\\ 
 -2 & 1 & 0\\
0 & 4 & -4 
\end{array}\right) = -8 - 16 - (-48) = -72 \neq 0", hence vectors are linearly independent.
  4. As in Task 2, solve corresponding linear system of equations: "\left(\begin{array}{ccc|c} 
 1 & 2 & 3 & 2\\ 
 0 & 2 & 3 & -1\\
 0 & 0 & 3 &3 
\end{array}\right)", from where "\lambda_1 = 3, \lambda_2 = -2," "\lambda_3 = 1", hence "\bold v = 3 \bold v_1 - 2 \bold v_2 + \bold v_3".